OpenStudy (anonymous):

Write the equation of the line that is tangent to the circle x2 + y2 = 100 at the point (8, 6).

OpenStudy (anonymous):

@homeschooler answer = 8x -6 y = 100 (equation of the line)

OpenStudy (anonymous):

its would be 8x -6 y = 100

OpenStudy (anonymous):

The question is multiple choice and neither of those are the answer. @cutelilcigarette @iamanexpert12

OpenStudy (anonymous):

what are the answer choices.

OpenStudy (anonymous):

A. y= -4/3x+50/3 B. y=3/4x-25/2 C. y=4/3x+25 D. y= -4x+50 @deshawn1

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Can you help? @deshawn1

OpenStudy (anonymous):

i can try

OpenStudy (anonymous):

Thanks! @deshawn1

OpenStudy (anonymous):

look ast this https://answers.yahoo.com/question/index?qid=20120331150722AAqmNcf

OpenStudy (anonymous):

it is not the same question but it should help